# Vizing’s Conjecture: A Density-Based Re-framing

### v1

@@ Abstract
We recast Vizing’s conjecture $\gamma(G\square H)\geq\gamma(G)\gamma(H)$ through a simple domination–density lens. Defining $\rho_{X}=\gamma(X)/|V(X)|$, the conjecture becomes $\rho_{G\square H}\geq\rho_{G}\rho_{H}$. Many classical estimates admit this normalised form, but if we select valid upper bounds $\rho_{G}\leq\tilde{\rho}_{G}$, $\rho_{H}\leq\tilde{\rho}_{H}$ and a product lower bound $\rho_{G\square H}\geq\tilde{\rho}_{G\square H}$, then the single test $\tilde{\rho}_{G\square H}\geq\tilde{\rho}_{G}\tilde{\rho}_{H}$ certifies Vizing’s. Instantiations give: (i) a bipartite imbalance criterion, where partition asymmetry together with a maximum–degree bound yields infinite nontrivial families; and (ii) a regular high–degree regime via the Arnautov–Payan bound, certifying all $k$-regular pairs for $k\geq 32$. The framework is modular: sharper family–specific bounds can be dropped in without altering the proof, expanding the certified regime while preserving the simple density form.
@@ 1 Domination Density Framework
Define domination densities:
@@
Then Vizing’s conjecture is equivalent to:
@@
Throughout, let $\Delta(X)$ and $\delta(X)$ denote the maximum and minimum degrees of a graph $X$.
@@ 2 Introduction
Here $\gamma(\cdot)$ denotes the domination number and $G\square H$ is the Cartesian product. The conjecture has withstood more than five decades of study: it is known to hold for various special classes (trees, cographs, interval graphs, among others), yet no general proof or counterexample has been found. Because domination numbers are notoriously difficult to compute, much of the progress has come from developing upper and lower bounds that can be combined to certify the inequality in restricted regimes.
The density framework introduced above provides a compact way to organise such bounds. This re-framing turns the search for sufficient conditions into a simple inequality between densities, highlighting the structural parameters that drive the conjecture.
In this short note we demonstrate the utility of this perspective through two case studies. First, we derive a bipartite imbalance criterion showing that sufficiently uneven bipartitions guarantee Vizing’s inequality, even for infinite nontrivial families. Second, we show how the framework recovers a regular high–degree regime, where the Arnautov–Payan bound certifies the conjecture for all kk–regular pairs once $k\geq 32$. These examples illustrate how the density lens not only unifies existing arguments but also provides a flexible template into which sharper bounds may be inserted to enlarge the certified regime.
@@ 3 Hybrid Sufficient Conditions and Practical Applications
In practical settings regarding Vizing’s conjecture, particularly for algorithmic searches or large-scale analysis, it is often advantageous to work with collections of known domination bounds tailored to specific graph families. That is, under such conditions, one can define sets of computationally light upper and lower domination density bounds, denoted $\mathrm{U}_{G}=\{\rho_{G1},\rho_{G2},\rho_{G3},\dots,\rho_{Gn}\}$ and $\mathrm{L}_{G}=\{\rho_{G1},\rho_{G2},\rho_{G3},\dots,\rho_{Gn}\}$, where each $\rho_{Gi}$ is a valid bound applicable to the structure of $G$.
By evaluating each bound in $\mathrm{U}_{G}$ for a given graph, one can select the tightest available estimate (e.g., the minimal upper bound), and use this to test whether sufficient conditions for Vizing’s Conjecture are satisfied. This hybrid approach—combining theoretical bounds with computational heuristics—may yield practical methods for certifying Vizing’s conjecture across broad graph classes without requiring explicit computation. This framework may also yield theoretical results restricting the parameter space in which a counterexample to the conjecture could possibly exist.
The framework defined explicitly is as follows: Let $G$ and $H$ be simple finite graphs, let $\mathrm{U}_{G}$ and $\mathrm{U}_{H}$ be the sets of upper density bounds with $\tilde{\rho}_{G}$ and $\tilde{\rho}_{H}$ as the minimum element of said sets so that $\rho_{G}\leq\tilde{\rho}_{G}$, and $\rho_{H}\leq\tilde{\rho}_{H}$:
$\tilde{\rho}_{G}:=\min\left\{\mathrm{U_{G}}\right\},\qquad\tilde{\rho}_{H}:=\min\left\{\mathrm{U_{H}}\right\},$ (3)
Similarly let $\mathrm{L_{G\square H}}$ be the set of lower density bounds applicable to the Cartesian product graph of $G$ and $H$ with $\tilde{\rho}_{G\square H}$ as the maximum element so that $\rho_{G\square H}\geq\tilde{\rho}_{G\square H}$:
$\tilde{\rho}_{G\square H}:=\max\left\{\mathrm{L_{G\square H}}\right\},$ (4)
@@
Extensions to this framework arise by incorporating bounds to curated sets associated with specific graph families. With development, this framework could become a powerful computational and theoretical tool for estimating domination numbers. Through the curation of diverse tight bounds, the theoretical parameter regime for potential contradictory graphs tightens, reducing the search space required to confirm Vizing’s conjecture. Such restrictions may aid in future theoretical proofs which could show transformations that expand the safe regime where Vizing’s conjecture is guaranteed.
4 Bipartite graphs: an imbalance sufficient condition
In this section we apply the framework above using simple domination bounds for bipartite graph pairs to define a small but non-trivial parameter regime for when bipartite graphs hold under Vizing’s conjecture.
Let $G$ and $H$ be connected bipartite graphs. Write the bipartition of $G$ as $V(G)=A_{G}\cup B_{G}$ with $A_{G}\cap B_{G}=\varnothing$, and set
$n_{1}(G):=|A_{G}|,\qquad n_{2}(G):=|B_{G}|,\qquad n_{1}(G)\leq n_{2}(G).$
Similarly, $V(H)=A_{H}\cup B_{H}$ with $n_{1}(H):=|A_{H}|$, $n_{2}(H):=|B_{H}|$, and $n_{1}(H)\leq n_{2}(H)$. Thus
$|V(G)|=n_{1}(G)+n_{2}(G),\qquad|V(H)|=n_{1}(H)+n_{2}(H),$
and
@@ 4 Bipartite graphs: an imbalance sufficient condition
For connected bipartite $X$, the smaller part dominates $X$, hence
$\gamma(X)\leq\min\{n_{1}(X),n_{2}(X)\}=n_{1}(X).$
Therefore the upper density bounds
$\tilde{\rho}_{G}:=\frac{n_{1}(G)}{n_{1}(G)+n_{2}(G)},\qquad\tilde{\rho}_{H}:=\frac{n_{1}(H)}{n_{1}(H)+n_{2}(H)}$
are valid for all connected bipartite $G,H$.
For any connected graph $X$ we also use the standard bound
$\gamma(X)\ \geq\ \frac{|V(X)|}{\Delta(X)+1},$
so for the Cartesian product we have $\tilde{\rho}_{G\square H}:=1/(\Delta(G)+\Delta(H)+1)$. Applying these bounds, the condition $\tilde{\rho}_{G\square H}\geq\tilde{\rho}_{G}\tilde{\rho}_{H}$ yields:
Lemma 1 (Imbalance–degree criterion).
Let $G,H$ be connected bipartite graphs with bipartition sizes $n_{1}(G)\leq n_{2}(G)$ and $n_{1}(H)\leq n_{2}(H)$. If
@@
equivalently
$\frac{n_{2}(G)}{n_{1}(G)}+\frac{n_{2}(H)}{n_{1}(H)}+\frac{n_{2}(G)}{n_{1}(G)}\frac{n_{2}(H)}{n_{1}(H)}\geq\Delta(G)+\Delta(H),$
then Vizing’s inequality holds for the pair $(G,H)$:
$\gamma(G\square H)\;\geq\;\gamma(G)\,\gamma(H).$
@@
Fix $\Delta(G)+\Delta(H)$ and $\frac{n_{2}(H)}{n_{1}(H)}$. The condition in Lemma 1 holds whenever
@@
In particular, $\frac{n_{2}(G)}{n_{1}(G)}\geq\Delta(G)+\Delta(H)$ or $\frac{n_{2}(H)}{n_{1}(H)}\geq\Delta(G)+\Delta(H)$ suffices, yielding infinite families via the constructions listed below.
Given the limitations of the parameter regime, we list constructions that yield arbitrarily large, structurally rich bipartite graphs which satisfy the sufficient condition. Note that this is easiest to satisfy when $G$ and $H$ are very large with sufficiently uneven bipartitions, i.e. when $n_{1}\ll n_{2}$. Constructions include:
  1. 1.
    One-sided blow ups via leaves and path tree-like structures. Given any initial bipartite core, the addition of one-sided leaves, longer pendant paths, as well as other structures which increase bipartition unevenness push a graph toward the successful parameter regime. Such operations can be repeated indefinitely such that the bipartition becomes sufficiently uneven, i.e. $n_{1}\ll n_{2}$.
  2. 2.
    Even-cycle grafting. Begin with a bipartite graph with an uneven partition. Select a vertex from the larger side and graft an even cycle $C_{2m}$ by identifying one of its vertices with the host. The cycle contributes evenly to both sides, so attaching it to the larger side preserves the imbalance while adding structural complexity (extra cycles, greater diameter, non-tree-like form). Repeating at multiple sites yields dense bipartite families rich in even cycles, yet still uneven, ensuring $\tilde{\rho}_{G\square H}\geq\tilde{\rho}_{G}\tilde{\rho}_{H}$ holds.
@@ 5 Regular graphs: a high–degree regime
Applying the same framework to regular graphs we first note that for a $k$-regular graph $X$ we have $\delta(X)=\Delta(X)=k$. We also recall that for Cartesian products,
@@
$\gamma(X)\ \leq\ \frac{H_{\delta(X)+1}}{\delta(X)+1}\,|V(X)|,\qquad H_{m}=\sum_{j=1}^{m}\frac{1}{j}.$

### v2

@@ Abstract
We present an equivalent form of Vizing’s conjecture $\gamma(G\square H)\geq\gamma(G)\gamma(H)$ using a simple domination–density lens. Defining $\rho_{G}=\gamma(G)/|V(G)|$, the conjecture becomes $\rho_{G\square H}\geq\rho_{G}\rho_{H}$, that is the dominating density of the Cartesian product is equal to or larger than the product of the domination densities of the original graphs. Selecting valid upper bounds such $\rho_{G}\leq\tilde{\rho}_{G}$, $\rho_{H}\leq\tilde{\rho}_{H}$ and a product lower bound such that $\rho_{G\square H}\geq\tilde{\rho}_{G\square H}$, then we note the simple test $\tilde{\rho}_{G\square H}\geq\tilde{\rho}_{G}\tilde{\rho}_{H}$ which certifies Vizing’s if holds. Examples of this give Vizing’s holds for: (i) bipartite graph pairs whose bi-partitions are sufficiently uneven compared to the graphs maximum degree, yielding infinite nontrivial graph pairs; and (ii) via the Arnautov–Payan bound for all $k$-regular pairs for $k\geq 32$. The framework utilised is modular and the bounds included quite general. The implementation of increasingly sharp family–specific bounds can be utilised to expand the certified parameter regime’s for when Vizing’s holds.
@@ 1 Domination Density Framework
Let $\Delta(G)$ and $\delta(G)$ denote the maximum and minimum degrees of a graph $G$. Define domination densities:
@@
Which implies Vizing’s conjecture is equivalent to:
@@ 2 Introduction
Here $\gamma(G)$ denotes the domination number and $G\square H$ is the Cartesian product. The conjecture is known to hold for various special classes (trees, cographs, interval graphs, among others), yet no general proof or counterexample has been found. Because domination numbers are difficult to compute, much progress has come from developing upper and lower bounds that can be combined to certify the inequality in restricted regimes.
We demonstrate the utility of this via two cases: First, we derive a bipartite regime showing that sufficiently uneven bi-partitions guarantee Vizing’s inequality for infinite nontrivial graph families. Second, we use the Arnautov–Payan bound to certify the conjecture for all kk–regular pairs $k\geq 32$. These illustrate how the density lens not only unifies existing arguments but also provides a flexible template into which sharper bounds may be inserted to enlarge the certified regime.
@@ 3 Hybrid Sufficient Conditions and Practical Applications
In practical settings regarding Vizing’s conjecture it is often advantageous to work with collections of known domination bounds tailored to specific graph families. That is, one can define sets of computationally light upper and lower domination density bounds which we denote $\mathrm{U}(G)=\{\rho_{G1},\rho_{G2},\rho_{G3},\dots,\rho_{Gn}\}$ and $\mathrm{L}_{G}=\{\rho_{G1},\rho_{G2},\rho_{G3},\dots,\rho_{Gn}\}$, where each $\rho_{Gi}$ is a valid lower of upper bound applicable to the structure of $G$.
By evaluating element in $\mathrm{U}(G)$ for a given graph, one can select the minimal upper bound and use this to test if sufficient conditions for Vizing’s Conjecture are satisfied. This hybrid approach combining theoretical bounds with computational heuristics may yield theoretical results restricting the parameter space in which a counterexample could possibly exist.
The framework is as follows: Let $G$ and $H$ be simple finite graphs. Let $\mathrm{U}(G)$ and $\mathrm{U}(H)$ be the sets of upper density bounds with $\tilde{\rho}_{G}$ and $\tilde{\rho}_{H}$ as the minimum element so that $\rho_{G}\leq\tilde{\rho}_{G}$, and $\rho_{H}\leq\tilde{\rho}_{H}$:
$\tilde{\rho}_{G}:=\min\left\{\mathrm{U}(G)\right\},\qquad\tilde{\rho}_{H}:=\min\left\{\mathrm{U}(H)\right\},$ (3)
Similarly let $\mathrm{L}(G\square H)$ be the set of lower density bounds applicable to the Cartesian product graph of $G$ and $H$ with $\tilde{\rho}_{G\square H}$ as the maximum element so that $\rho_{G\square H}\geq\tilde{\rho}_{G\square H}$:
$\tilde{\rho}_{G\square H}:=\max\left\{\mathrm{L}(G\square H)\right\},$ (4)
@@
Corollary 1 (Order–invariance of domination density).
Domination densities are invariant under graph order. Consequently, when averaging over the induced sub-graphs of $G\square H$, the mean domination density satisfies
$\overline{\rho}_{\,G\square H}\;\geq\;\rho_{G}\,\rho_{H}.$
Proof.
By definition $\rho_{G}=\gamma(G)/|V(G)|$ and $\rho_{H}=\gamma(H)/|V(H)|$, both of which are ratios independent of scaling in $|V(G)|,|V(H)|$. Thus, densities are order invariant. For each induced sub-graph of $G\square H$, the same density property holds, and by averaging over all such sub-graphs the resulting mean density cannot be less than the product $\rho_{G}\rho_{H}$, which is the lower threshold enforced by Vizing’s inequality. ∎
Through the curation of diverse tight bounds, the theoretical parameter regime for potential contradictory graphs tightens, reducing the search space to confirm Vizing’s conjecture. Such restrictions may aid in future theoretical proofs which could show transformations that expand the safe regime where Vizing’s conjecture is guaranteed.
4 Bipartite Graphs: Bi-partition imbalance regime
In the following section we combine the domination density framework with simple domination bounds applicable to bipartite graphs to derive a bi-partition imbalance regime for when pairs of bipartite graphs satisfy Vizing’s conjecture. Let $G$ and $H$ be connected bipartite graphs. Write the bi-partition of $G$ as $V(G)=A_{G}\cup B_{G}$ with $A_{G}\cap B_{G}=\emptyset$, and set $n_{1}(G):=|A_{G}|$, $n_{2}(G):=|B_{G}|$, and $n_{1}(G)\leq n_{2}(G).$ Similarly for $H$, $V(H)=A_{H}\cup B_{H}$ with $n_{1}(H):=|A_{H}|$, $n_{2}(H):=|B_{H}|$, and $n_{1}(H)\leq n_{2}(H)$. Thus
$|V(G)|=n_{1}(G)+n_{2}(G)\qquad\text{and}\qquad|V(H)|=n_{1}(H)+n_{2}(H)$
where
@@ 4 Bipartite Graphs: Bi-partition imbalance regime
We note that for any connected bipartite graph $G$ the nodes of the smaller bi-partition always dominate $G$, hence
$\gamma(G)\leq n_{1}(G)$
which allows us to define the following bounds which are valid for all connected bipartite graphs
$\rho_{G}\leq\frac{n_{1}(G)}{n_{1}(G)+n_{2}(G)}\qquad\text{and}\qquad\rho_{H}\leq\frac{n_{1}(H)}{n_{1}(H)+n_{2}(H)}.$
Let $\Delta(G)$ and $\delta(G)$ denote the maximum and minimum degrees of a graph $G$. We then note that any connected graph $G$ satisfies the standard bound
$\gamma(G)\ \geq\ \frac{|V(G)|}{\Delta(G)+1},$
so for the Cartesian product, given $\Delta(G\square H)=\Delta(G)+\Delta(H)$, we have
$\rho_{G\square H}\geq\frac{1}{\Delta(G)+\Delta(H)+1}$
Applying these bounds, the condition $\rho_{G\square H}\geq\rho_{G}\rho_{H}$ yields:
Theorem 1 (Imbalance–degree criterion).
Let $G,H$ be connected bipartite graphs with bi-partition sizes $n_{1}(G)\leq n_{2}(G)$ and $n_{1}(H)\leq n_{2}(H)$. If
@@
then Vizing’s inequality holds for the pair $(G,H)$, i.e. $\gamma(G\square H)\geq\gamma(G)\gamma(H)$.
@@
Fix $\Delta(G)+\Delta(H)$ and $\frac{n_{2}(H)}{n_{1}(H)}$. The condition in Theorem 1 holds whenever
@@
In particular, $\frac{n_{2}(G)}{n_{1}(G)}\geq\Delta(G)+\Delta(H)$ or $\frac{n_{2}(H)}{n_{1}(H)}\geq\Delta(G)+\Delta(H)$ suffices.
@@ 5 Regular graphs: a high–degree regime
Applying the same framework to regular graphs we first note that for a $k$-regular graph $G$ we have $\delta(G)=\Delta(G)=k$. We also recall that for Cartesian products,
@@
$\gamma(G)\ \leq\ \frac{H_{\delta(G)+1}}{\delta(G)+1}\,|V(G)|,\qquad H_{G}=\sum_{j=1}^{G}\frac{1}{j}.$